# Computational electro-optic frequency comb spectroscopy

### v1

@@ 1 Introduction
In contrast to other comb platforms, EO-OFCs offer easy implementation, robustness, and a flexibility that enables line spacing tuning across several orders of magnitude (from megahertz to tens of gigahertz). This comb category is also well suited for dual-comb spectroscopy (DCS) [10], as the generation of two EO-OFCs from a single cw laser inherently ensures high mutual coherence [11, 12]. A limitation of EO combs is their modest number of spectral lines, typically less than one hundred. The comb bandwidth, however, can be dramatically increased by non-linear broadening techniques [11, 13, 14] or by placing the EO modulator within a cavity [15, 16], albeit at the expense of reduced flexibility. EO combs have been successfully employed in a wide range of applications, including molecular spectroscopy [11, 17], distance measurement [18, 19], telecommunications [20], and, more recently, distributed fiber sensing [21, 22]. Beyond these established applications, the particular characteristics of EO combs make them especially well-suited for new computational approaches.
@@
Figure 1: Operational principle. A comb generator, driven by a dynamically reconfigurable signal and fed by a cw laser, generates a set of known comb spectra for interrogating a spectroscopic sample. In each reconfiguration, the power is redistributed among the comb lines but remains constant, as illustrated in the bottom-left inset (const denotes a constant value). In contrast, the power transmitted through (or reflected by) the sample varies sequentially based on the overlap between its spectral response and each generated comb. The sample’s response is then reconstructed from the measured power values by solving a linear system $\mathbf{Ax}=\mathbf{b}$, where $\mathbf{A}$ is the spectrum matrix, and $\mathbf{x}$ and $\mathbf{b}$ represent the spectral response and the measured powers, respectively (see bottom-right inset). In practice, this linear system is ill-posed, and its solution requires formulating and solving an optimization problem.
2 Operation principle
2.1 Driving signals
@@
For the results presented in this paper, it suffices to consider up to two harmonics. The more complex case of h-tone modulation is treated without a priori assumptions in Appendix A.
2.1.1 Single-tone RF signal
@@
where $\left\{J_{k}\right\}$ are Bessel functions of the first kind. For simplicity, we have omitted the subscript $l$ in the expression for $\Phi_{i}\left(t\right)$ and, without loss of generality, assumed that $\varphi_{i}=0$ $\forall$ ii. Considering $J_{-k}=\left(-1\right)^{k}J_{k}\left(k>0\right)$, the expression for the output field can be written as:
@@
The Fourier transform of this expression reveals a well-known result: the output power spectrum consists of lines evenly spaced by $f_{s}$ and symmetric about $\nu_{0}$. The comb’s shape and total number of lines $N$ (typically defined relative to a minimum power threshold) depend on the modulation index $\beta_{i}$, which can be adjusted by varying the applied peak voltage $u_{i}$. As a result, a sequence of comb spectra can be generated by driving a PM with a sinusoidal signal of frequency $f_{s}$ and an amplitude modulated in a staircase pattern, as we will show in Sec. IIB. In practice, the maximum $\beta_{i}$ is limited by the RF power handling capability of the PM. Figure 2(a) illustrates an example of a comb generated by applying a single-tone signal with a modulation index of $2\pi$.
2.1.2 Two-tone RF signal
@@
where we assume that $\varphi_{1i}=0$, allowing $\varphi_{2i}$ to be interpreted as the relative phase between the two tones comprising the voltage signal. For the above multi-harmonic phase, it remains possible to express $e^{j\Phi_{i}\left(t\right)}$ in Eq. (2) as a Jacobi-Anger expansion, but using two-variable, one-parameter generalized Bessel functions (GBFs) [34] as follows:
@@ 2.2 Reconstruction algorithm
$\mathbf{Ax}=\mathbf{b}.\\$ (10)
In this equation, $\mathbf{A}$ is a matrix formed by consecutive probe combs arranged as rows. The element $A_{mn}$ is the spectral power of the n-th comb line for the m-th input spectrum. The vector $\mathbf{x}$ represents the spectral response at the frequency of each comb line, while $\mathbf{b}$ denotes the sequence of powers measured during the data acquisition process. The limits $\nu_{1}$ and $\nu_{2}$ correspond to the frequencies of the first and last comb lines, respectively. From now on, $\mathbf{A}$ will always be considered a square matrix (that is, $\mathbf{x}$ and $\mathbf{b}$ will have the same length) to simplify the mathematical treatment.
The most straightforward way to solve the linear system described above is by multiplying the inverse of $\mathbf{A}$ by $\mathbf{b}$. However, as in many other applications, directly computing the inverse matrix leads to numerical errors or is simply not possible. To clarify this point, it is useful to employ the singular value decomposition (SVD) of $\mathbf{A}$, which allows us to express this matrix as $\mathbf{A}=\mathbf{U}\mathbf{S}\mathbf{V}^{T}$, where $\mathbf{U}$ and $\mathbf{V}$ are orthogonal matrices, $T$ denotes the transpose, and $\mathbf{S}$ is a diagonal matrix with positive real elements $S_{nn}=s_{n}$ (the singular values of $\mathbf{A}$). By convention, the values of $s_{n}$ are arranged in decreasing order. In our approach, the comb spectra generate a system where the ratio between the largest and smallest singular values is extremely high, making direct inversion of $\mathbf{A}$ impractical. This issue can be addressed using a ”truncated” SVD, where all diagonal elements below a given threshold $s_{th}$ are intentionally set to zero [35]. This operation facilitates the calculation of the (Moore-Penrose) ”pseudo-inverse” matrix, given by $\mathbf{A}^{\dagger}=\mathbf{V}\mathbf{S}^{\dagger}\mathbf{U}^{T}$, where $\mathbf{S}^{\dagger}$ contains the reciprocals of nonzero elements in the truncated version of $\mathbf{S}$. The reconstructed sample response is then calculated as $\mathbf{x}=\mathbf{A}^{\dagger}\mathbf{b}$. It can be proven that the solution $\mathbf{x}$ obtained in this way minimizes the Euclidean norm $\mathbf{Ax}-\mathbf{b}||_{2}$ [36].
@@
To determine the threshold $s_{th}$ of the truncated SVD, we plot the singular values of $\mathbf{A}$ as a function of their diagonal index on a logarithmic scale. As expected, they follow a decreasing smooth slope, which becomes significantly more abrupt from a certain element of $\mathbf{S}$. Setting all subsequent values equal to zero provides a matrix that minimizes the root mean square error (RMSE) of the reconstruction. Once the threshold parameter is optimized, and as long as the set of programmed combs remains identical, the truncated SVD can be applied to other spectral responses. More details about this procedure can be found in Sec. S1 of the Supplementary Material. The blue squares in Fig. 3(b) are the values calculated from the Moore-Penrose matrix, while the solid line represents the theoretical sample’s response. For these results, RMSE =0.018=0.018 when $s_{th}=0.01$. As can be observed, the reconstructed values accurately reproduce the central peak of the filter response, as well as two sidelobes on the left and right of the central frequency.
@@
The next example is a spectral response resulting from the overlap of two Gaussian absorption curves, one blue-shifted (centered at $6$ GHz) and the other red-shifted (centered at 2-2 GHz) relative to the laser frequency. Each curve has a different height and full width at half maximum (FWHM) (approximately $8$ GHz and $4$ GHz, respectively). In this case, using a driving signal like that in Fig. 3(a) leads to a numerical solution of Eq. (10) that has no physical meaning, since the reconstruction is always symmetric with respect to the laser frequency. This issue can be addressed by applying a two-harmonic signal to the PM to generate asymmetric probe combs (see Fig. 2). As in the previous example, we set the parameter $\beta_{1}/\pi$ (for the frequency $f_{s}$) to take values from 0 to a maximum level of $4$. The values of $\beta_{2}/\pi$ for the $2f_{s}$ signal are chosen to be identical to $\beta_{1}/\pi$ but $10$ times weaker, thus avoiding the practical need to amplify this additional component. We consider $f_{s}=1$ GHz and $M=41$. The specific values of $\beta_{2}$ and the relative phase $\varphi_{2}$ for each PM reconfiguration are selected as follows. We define a vector of values for $\beta_{2}/\pi$, which, as explained above, is formed by a set of $M$ equidistant levels from 0 to $0.4$. The phase $\varphi_{2}$ forms another vector of the same length as $\beta_{2}/\pi$ constructed by alternating the values 0 and $\pi$. Next, we randomly permute the elements of both sequences and use the resulting vectors to successively select the parameters of the $2f_{s}$ component at each measurement step. Introducing such a randomness into Eq. (10) facilitates the creation of dissimilar probe spectra [39]. However, depending on the available electronic equipment, the choice of the driving signal can be simplified, as will be discussed in Sec. III. A sketch of the optical phase generated for three measurement steps is shown in Fig. 3(c). As in the previous example, we add Gaussian noise to the powers $\left\{b_{i}\right\}(\sigma_{n}=0.003)$ and resolve Eq. (10) using a truncated SVD of $\textbf{A}(s_{th}=0.05)$. Fig. 3(d) shows that the recovered $x(\nu)$ fits the reference curve accurately (RMSE =0.016=0.016).
3 Experimental results
@@
Figure 5: Experimental results for the transmission $x(\nu)$ of a tunable bandpass filter centered at the laser frequency. (a) Series of measured powers $b_{i}$ as a function of the modulation index $\beta/\pi$. (b) Reconstructed values of $x(\nu)$ (blue squares) together with a reference curve (red line) obtained by fitting dual-comb measurements (green triangles) with a Gaussian function. The RMSE of the reconstruction is $0.02$.
@@
In the previous result, we have assumed precise alignment between the laser frequency and the center of the filter transmission, which is achievable because of the fine-wavelength tuning of our laser. However, this requirement can be relaxed by employing a two-tone driving signal. With the electronic equipment available in our lab, we can accurately control the relative phase $\varphi_{2}$ between an amplified AM signal ($f_{s}=500$ MHz) and a sinusoid ($2f_{s}=1$ GHz) using an electrical phase shifter operating at $500$ MHz (JSPHS-661+ evaluation board, Mini-Circuits). We consider only two values of $\varphi_{2}$~(0 and $\pi$), as in the spectra shown in Fig. 2. In this case, the PM driven by the highest RF power provides a maximum of 14 comb lines within 20 dB, requiring us to reduce the filter’s FWHM to the minimum attainable value (around $4$ GHz). Despite all these limitations, this second experiment allows us to reconstruct the filter response when there is a small relative shift between its center and the laser frequency. The voltage amplitude of the $500$ MHz signal is linearly varied (from $1$ to $7.6$ V). The half-wave voltage at this frequency is $5$ V. The power of the $1$ GHz sinusoid (coming from the second output of our signal generator) is set to a constant value that does not require amplification ($1$ V of peak amplitude). We consider $51$ reconfigurations, $26$ of them with $\varphi_{2}=0$ (including a measurement with the PM in the off-state) and the remaining $25$ with $\varphi_{2}=\pi$. This driving scheme can be implemented by sending two consecutive modulated signals with different values of $\varphi_{2}$. The numerically computed comb spectra are shown in Sec. S3 of the Supplementary Material. The laser frequency $\nu_{0}$ is shifted $300$ MHz ($0.6~f_{s}$). The measured powers $\{b_{i}\}$ (Fig. 6(a)) are the result of averaging $5\times 10^{4}$ points ($\sigma_{n}=0.003$). These measurements demonstrate that the two values of $\varphi_{2}$ (corresponding, respectively, to the dark and light green points in Fig. 6(a)) make it possible to distinguish between the left side and the right side of the filter response with respect to $\nu_{0}$. The reconstructed transmission over $12$ GHz is shown in Fig. 6(b), along with the reference curve obtained by interpolating the points measured by DCI using a line spacing of $1$ GHz (RMSE =0.05=0.05). A zoom-in view of the maximum transmission area (inset in Fig. 6(b)) shows again a good fit between both curves, clearly revealing the frequency shift with respect to the laser frequency. Experimentally, one can observe that if $\nu_{0}$ shifts toward the center of the transmission curve, the dark and light green points in Fig. 6 overlap closely, resulting in a single curve, as expected.
@@
Figure 6: Experimental results for the transmission $x(\nu)$ of a tunable bandpass filter shifted with respect to the laser frequency. (a) Series of measured powers $b_{i}$ as a function of the modulation index $\beta_{1}/\pi$ for two values (0 and $180\textdegree$) of the relative shift between the signals at $f_{s}$ and $2f_{s}$. (b) Reconstructed values of $x(\nu)$ (blue squares) together with a reference curve (red line) obtained by fitting dual-comb measurements (green triangles) with a Gaussian function. The RMSE of the reconstruction is $0.05$. The inset shows a zoomed-in view of the peak maximum.
@@
As a final example, we consider a molecular transition line as a spectroscopic sample. The transmission $T$ for such a fingerprint signature appears as a dip on a constant background at a well-defined frequency. In contrast, the power of the lines of our probe combs eventually tends to zero as they move far enough from the central frequency. Consequently, in this case, the EO combs generated by our system are not suitable sampling functions. This issue can be addressed by removing the background (signal in the absence of transition), which leads to the reconstruction of a function with a behavior like that in the previous examples. A similar strategy is employed in compressive dual-comb spectroscopy to ensure sparsity in the spectral responses [41]. For this purpose, we take an additional power measurement $b_{0}$ far from the center of the molecular transition. The curve $x^{\prime}=1-T$ is then reconstructed using the linear system of Eq. (10), but with a modified sequence of powers $\{b^{\prime}_{i}\}$, where $b^{\prime}_{i}=1-b_{i}/b_{0}$ (see Appendix C).
@@
In our experiment, we use a gas cell with an absorption length of $5.5$ cm that contains $\text{H}^{\text{13}}\text{CN}$ at $3.3$ kPa ($25$ Torr), with a $10\%$ uncertainty in pressure (Wavelength References). For such a low pressure, a good reconstruction of the absorption line under consideration (P4 at $1545.23$ nm) requires a sampling comb with a line spacing of (at most) $500$ MHz and an optical bandwidth extending on the order of $10$ GHz. In our setup, a single phase modulator cannot meet the above requirements, as the maximum achievable modulation index (approximately $1.7\pi$) is insufficient to generate the required number of comb lines. To overcome this limitation, we implement a comb generator consisting of two cascaded PMs. When driven by in-phase sinusoidal signals of identical frequency, both PMs behave as a single modulator with $\beta=\beta_{1}+\beta_{2}$, where $\beta_{1}$ and $\beta_{2}$ are the modulation indices of each PM [42]. For this experiment, we align the laser with the transition frequency and use a single-tone AM signal (like in Fig. 3(a)). This driving signal is split into two arms and our phase shifter is inserted in one of them to ensure precise synchronization between both PMs. In each path, identical RF amplifiers are inserted (ABP0-0300-01-3830, Wenteq Microwave). The half-wave voltage at $500$ MHz for the second PM is $4.9$ V. The values of $\beta_{1}/\pi$ and $\beta_{2}/\pi$ undergo an approximately linear variation from $0.88$ to $1.64$ and from $0.73$ to $1.68$, respectively. With a sufficiently high modulation index, a single PM could be employed to generate the combs, while the phase shifter would assist in producing a two-tone modulation, thereby eliminating the need to align the laser with the center of the absorption line (see Sec. II). As in the previous experiment, we consider $51$ reconfigurations (including an off-state measurement), so the acquisition time is $10$ ms. Each $b^{\prime}_{i}$ is the result of averaging $5\times 10^{4}$ points ($\sigma_{n}=0.001$). By solving the corresponding linear system ($s_{th}=0.03$), we reconstruct $x^{\prime}(v)$. As a reference, we used the P4 line shape for the HCN at $25$ Torr, derived from the NIST data. The reconstructed points spanning over $13$ GHz, shown in Fig. 7(b), closely match the reference values (RMSE =0.015=0.015).
@@
Figure 7: Experimental results for the spectrum $x^{\prime}=1-T$, where T is the transmission of a $5.5$ cm-long HCN gas cell at $25$ Torr, centered at $1545.23$ nm. (a) Series of the measured powers $b_{i}$ as a function of the modulation index $\beta/\pi$. (b) Reconstructed spectral values of $x^{\prime}$, (blue squares) and the reference curve (red line) obtained from NIST data. The RMSE of the reconstruction is $0.015$.
Table 1: Summary of parameter values from simulations and experiments

### v2

@@ I Introduction
In contrast to other comb platforms, EO-OFCs offer easy implementation, robustness, and a flexibility that enables line spacing tuning across several orders of magnitude (from megahertz to tens of gigahertz). This comb category is also well suited for dual-comb spectroscopy (DCS) [10], as the generation of two EO-OFCs from a single cw laser inherently ensures high mutual coherence [11, 12]. A limitation of EO combs is their modest number of spectral lines, typically less than one hundred. The comb bandwidth, however, can be dramatically increased by non-linear broadening techniques [11, 13, 14] or by placing the EO modulator within a cavity [15, 16], albeit at the expense of reduced flexibility. EO combs have been successfully employed in a wide range of applications, including molecular spectroscopy [11, 17], distance measurement [18, 19], telecommunications [20], and, more recently, distributed fiber sensing [21, 22]. Beyond these established applications, the particular characteristics of EO combs make them especially appropriate for new computational approaches.
@@
Figure 1: Operational principle. A comb generator, driven by a dynamically reconfigurable signal and fed by a cw laser, generates a set of known comb spectra for interrogating a spectroscopic sample. In each reconfiguration, the power is redistributed among the comb lines but remains constant, as illustrated in the bottom-left inset (const denotes a constant value). In contrast, the power transmitted through (or reflected by) the sample varies sequentially based on the overlap between its spectral response and each generated comb. The sample’s response is then reconstructed from the measured power values by solving a linear system $\textbf{Ax}=\textbf{b}$, where A is the spectrum matrix, and x and b represent the spectral response and the measured powers, respectively (see bottom-right inset). In practice, this linear system is ill-posed, and its solution requires formulating and solving an optimization problem.
II Operation principle
II-A Driving signals
@@
For the results presented in this paper, it suffices to consider up to two harmonics. The more complex case of h-tone modulation is treated without any simplifying assumptions in Appendix A.
II-A1 Single-tone RF signal
@@
where $\left\{J_{k}\right\}$ are Bessel functions of the first kind. For simplicity, we have omitted the subscript $l$ in the expression for $\Phi_{i}\left(t\right)$ given by Eq. (3) and, without loss of generality, assumed that $\varphi_{i}=0$ $\forall$ ii. Taking into account $J_{-k}=\left(-1\right)^{k}J_{k}\left(k>0\right)$, the output field can be written as:
@@
The Fourier transform of this equation reveals a well-known result: the output power spectrum consists of lines evenly spaced by $f_{s}$ and symmetric about $\nu_{0}$. The comb’s shape and total number of lines $N$ (typically defined relative to a minimum power threshold) depend on the modulation index $\beta_{i}$, which can be adjusted by varying the applied peak voltage $u_{i}$. As a result, a sequence of comb spectra can be generated by driving a PM with a sinusoidal signal of frequency $f_{s}$ and an amplitude modulated in a staircase pattern, as we will show in Sec. IIB. In practice, the maximum $\beta_{i}$ is limited by the RF power handling capability of the PM. Figure 2(a) illustrates an example of a comb generated by applying a single-tone signal with a modulation index of $2\pi$.
II-A2 Two-tone RF signal
@@
where we assume that $\varphi_{1i}=0$, allowing $\varphi_{2i}$ to be interpreted as the relative phase between the two tones comprising the voltage signal. For the above two-harmonic phase, it remains possible to express $e^{j\Phi_{i}\left(t\right)}$ in Eq. (2) as a Jacobi-Anger expansion, but using two-variable, one-parameter generalized Bessel functions (GBFs) [34] as follows:
@@ II-B Reconstruction algorithm
$\textbf{Ax}=\textbf{b}.\\$ (10)
In this equation, A is a matrix formed by consecutive probe combs arranged as rows. The element $A_{mn}$ is the spectral power of the n-th comb line for the m-th input spectrum. The vector x represents the spectral response at the frequency of each comb line, while b denotes the sequence of powers measured during the data acquisition process. The limits $\nu_{1}$ and $\nu_{2}$ correspond to the frequencies of the first and last comb lines, respectively. From now on, A will always be considered a square matrix (that is, x and b will have the same length) to simplify the mathematical treatment.
The most straightforward way to solve the linear system described above is by multiplying the inverse of A by b. However, as in many other applications, directly computing the inverse matrix leads to numerical errors or is simply not possible. To clarify this point, it is useful to employ the singular value decomposition (SVD) of A, which allows us to express this matrix as $\textbf{A}=\textbf{U}\textbf{S}\textbf{V}^{T}$, where U and V are orthogonal matrices, $T$ denotes the transpose, and S is a diagonal matrix with positive real elements $S_{nn}=s_{n}$ (the singular values of A). By convention, the values of $s_{n}$ are arranged in decreasing order. In our approach, the comb spectra generate a system where the ratio between the largest and smallest singular values is extremely high, making direct inversion of A impractical. This issue can be addressed using a ”truncated” SVD, where all diagonal elements below a given threshold $s_{th}$ are intentionally set to zero [35]. This operation facilitates the calculation of the (Moore-Penrose) ”pseudo-inverse” matrix, given by $\textbf{A}^{\dagger}=\textbf{V}\textbf{S}^{\dagger}\textbf{U}^{T}$, where $\textbf{S}^{\dagger}$ contains the reciprocals of nonzero elements in the truncated version of S. The reconstructed sample response is then calculated as $\textbf{x}=\textbf{A}^{\dagger}\textbf{b}$. It can be proven that the solution x obtained in this way minimizes the Euclidean norm $\textbf{Ax}-\textbf{b}||_{2}$ [36].
@@
To determine the threshold $s_{th}$ of the truncated SVD, we plot the singular values of A as a function of their diagonal index on a logarithmic scale. As expected, they follow a decreasing smooth slope, which becomes significantly more abrupt from a certain element of S. Setting all subsequent values equal to zero provides a matrix that minimizes the root mean square error (RMSE) of the reconstruction. Once the threshold parameter is optimized, and as long as the set of programmed combs remains identical, the truncated SVD can be applied to other spectral responses. More details about this procedure can be found in Sec. S1 of the Supplementary Material. The blue squares in Fig. 3(b) are the values calculated from the Moore-Penrose matrix, while the solid line represents the theoretical sample’s response. For these results, RMSE =0.018=0.018 when $s_{th}=0.01$. As can be observed, the reconstructed values accurately reproduce the central peak of the filter response, as well as two sidelobes on the left and right of the central frequency.
The next example is a spectral response resulting from the overlap of two Gaussian absorption curves, one blue-shifted (centered at $6$ GHz) and the other red-shifted (centered at 2-2 GHz) relative to the laser frequency. Each curve has a different height and full width at half maximum (FWHM) (approximately $8$ GHz and $4$ GHz, respectively). In this case, using a driving signal like that in Fig. 3(a) leads to a numerical solution of Eq. (10) that has no physical meaning, since the reconstruction is always symmetric with respect to the laser frequency. This issue can be addressed by applying a two-harmonic signal to the PM to generate asymmetric probe combs (see Fig. 2). As in the previous example, we set the parameter $\beta_{1}/\pi$ (for the frequency $f_{s}$) to take values from 0 to a maximum level of $4$. The values of $\beta_{2}/\pi$ for the $2f_{s}$ signal are chosen to be identical to $\beta_{1}/\pi$ but $10$ times weaker, thus avoiding the practical need to amplify this additional component. We consider $f_{s}=1$ GHz and $M=41$. The specific values of $\beta_{2}$ and the relative phase $\varphi_{2}$ for each PM reconfiguration are selected as follows. We define a vector of values for $\beta_{2}/\pi$, which, as explained above, is formed by a set of $M$ equidistant levels from 0 to $0.4$. The phase $\varphi_{2}$ forms another vector of the same length as $\beta_{2}/\pi$ constructed by alternating the values 0 and $\pi$. Next, we randomly permute the elements of both sequences and use the resulting vectors to successively select the parameters of the $2f_{s}$ component at each measurement step. Introducing such a randomness into Eq. (10) facilitates the creation of dissimilar probe spectra [39]. However, depending on the available electronic equipment, the choice of the driving signal can be simplified, as will be discussed in Sec. III. A sketch of the optical phase generated for three measurement steps is shown in Fig. 3(c). As in the previous example, we add Gaussian noise to the powers $\left\{b_{i}\right\}(\sigma_{n}=0.003)$ and solve Eq. (10) using a truncated SVD of $\textbf{A}(s_{th}=0.05)$. Fig. 3(d) shows that the recovered $x(\nu)$ fits the reference curve accurately (RMSE =0.016=0.016).
@@
III EXPERIMENTAL RESULTS
@@
Figure 5: Experimental results for the transmission $x(\nu)$ of a tunable bandpass filter centered at the laser frequency. (a) Series of measured powers $b_{i}$ as a function of the modulation index $\beta/\pi$. (b) Reconstructed values of $x(\nu)$ (blue squares) together with a reference curve (red line) obtained by fitting dual-comb measurements (green triangles) with a Gaussian function. The RMSE of the reconstruction is $0.018$.
@@
In the previous result, we have assumed precise alignment between the laser frequency and the center of the filter transmission, which is achievable because of the fine-wavelength tuning of our laser. However, this requirement can be relaxed by employing a two-tone driving signal. With the electronic equipment available in our lab, we can accurately control the relative phase $\varphi_{2}$ between an amplified AM signal ($f_{s}=500$ MHz) and a sinusoid ($2f_{s}=1$ GHz) using an electrical phase shifter operating at $500$ MHz (JSPHS-661+ evaluation board, Mini-Circuits). We consider only two values of $\varphi_{2}$~(0 and $\pi$), as in the spectra shown in Fig. 2. In this case, the PM driven by the highest RF power provides a maximum of 14 comb lines within 20 dB, requiring us to reduce the filter’s FWHM to the minimum attainable value (around $4$ GHz). Despite all these limitations, this second experiment allows us to reconstruct the filter response when there is a small relative shift between its center and the laser frequency. The voltage amplitude of the $500$ MHz signal is linearly varied (from $0.9$ to $7.6$ V). The half-wave voltage at this frequency is $5$ V. The power of the $1$ GHz sinusoid (coming from the second output of our signal generator) is set to a constant value that does not require amplification ($1$ V of peak amplitude). We consider $51$ reconfigurations, $26$ of them with $\varphi_{2}=0$ (including a measurement with the PM in the off-state) and the remaining $25$ with $\varphi_{2}=\pi$. This driving scheme can be implemented by sending two consecutive modulated signals, each with a different value of $\varphi_{2}$. The numerically computed comb spectra are shown in Sec. S3 of the Supplementary Material. The laser frequency $\nu_{0}$ is shifted $300$ MHz ($0.6f_{s}$). The measured powers $\{b_{i}\}$ (Fig. 6(a)) are the result of averaging $5\times 10^{4}$ points ($\sigma_{n}=0.003$). These measurements demonstrate that the two values of $\varphi_{2}$ (corresponding, respectively, to the dark and light green points in Fig. 6(a)) make it possible to distinguish between the left side and the right side of the filter response with respect to $\nu_{0}$. The reconstructed transmission over $12$ GHz ($s_{th}=0.05$) is shown in Fig. 6(b), along with the reference curve obtained by interpolating the points measured by DCI using a line spacing of $1$ GHz (RMSE =0.070=0.070). A zoom-in view of the maximum transmission area (inset in Fig. 6(b)) shows again a good fit between both curves, clearly revealing the frequency shift with respect to the laser frequency. Experimentally, one can observe that if $\nu_{0}$ shifts toward the center of the transmission curve, the dark and light green points in Fig. 6 overlap closely, resulting in a single curve, as expected.
@@
Figure 6: Experimental results for the transmission $x(\nu)$ of a tunable bandpass filter shifted with respect to the laser frequency. (a) Series of measured powers $b_{i}$ as a function of the modulation index $\beta_{1}/\pi$ for two values (0 and $180\textdegree$) of the relative shift between the signals at $f_{s}$ and $2f_{s}$. (b) Reconstructed values of $x(\nu)$ (blue squares) together with a reference curve (red line) obtained by fitting dual-comb measurements (green triangles) with a Gaussian function. The RMSE of the reconstruction is $0.07$. The inset shows a zoomed-in view of the peak maximum.
@@
As a final example, we consider a molecular transition line as a spectroscopic sample. The transmission $T$ for such a fingerprint signature appears as a dip on a constant background at a well-defined frequency. In contrast, the power of the lines of our probe combs eventually tends to zero as they move far enough from the central frequency. Consequently, in this case, the EO combs generated by our system are not suitable sampling functions. This issue can be addressed by removing the background (i.e., the signal in the absence of transition), which leads to the reconstruction of a function with a behavior like that in the previous examples. A similar strategy is employed in compressive dual-comb spectroscopy to ensure sparsity in the spectral responses [41]. For this purpose, we take an additional power measurement $b_{0}$ far from the center of the molecular transition. The curve $x^{\prime}=1-T$ is then reconstructed using the linear system of Eq. (10), but with a modified sequence of powers $\{b^{\prime}_{i}\}$, where $b^{\prime}_{i}=1-b_{i}/b_{0}$ (see Appendix C).
@@
In our experiment, we use a gas cell with an absorption length of $5.5$ cm that contains $\text{H}^{\text{13}}\text{CN}$ at $3.3$ kPa ($25$ Torr), with a $10\%$ uncertainty in pressure (Wavelength References). For such a low pressure, a good reconstruction of the absorption line under consideration (P4 at $1545.23$ nm) requires a sampling comb with a line spacing of (at most) $500$ MHz and an optical bandwidth extending on the order of $10$ GHz. In our setup, a single phase modulator cannot meet the above requirements, as the maximum achievable modulation index (approximately $1.7\pi$) is insufficient to generate the required number of comb lines. To overcome this limitation, we implement a comb generator consisting of two cascaded PMs. When driven by in-phase sinusoidal signals of identical frequency, both PMs behave as a single modulator with $\beta=\beta_{1}+\beta_{2}$, where $\beta_{1}$ and $\beta_{2}$ are the modulation indices of each PM [42]. For this experiment, we align the laser with the transition frequency and use a single-tone AM signal (like in Fig. 3(a)). This driving signal is split into two arms and our phase shifter is inserted in one of them to ensure precise synchronization between both PMs. In each path, identical RF amplifiers are inserted (ABP0-0300-01-3830, Wenteq Microwave). The half-wave voltage at $500$ MHz for the second PM is $4.9$ V. The values of $\beta_{1}/\pi$ and $\beta_{2}/\pi$ undergo an approximately linear variation from $0.18$ to $1.64$ and from $0.15$ to $1.68$, respectively. With a sufficiently high modulation index, a single PM could be employed to generate the combs, while the phase shifter would assist in producing a two-tone modulation, thereby eliminating the need to align the laser with the center of the absorption line (see Sec. II). As in the previous experiment, we consider $51$ reconfigurations (including an off-state measurement), so the acquisition time is $10$ ms, see Fig. 7(a). Each $b^{\prime}_{i}$ is the result of averaging $5\times 10^{4}$ points ($\sigma_{n}=0.001$). By solving the corresponding linear system ($s_{th}=0.03$), we reconstruct $x^{\prime}(v)$. As a reference, we used the P4 line shape for the HCN at $25$ Torr, derived from the NIST data. The reconstructed points spanning over $13$ GHz, shown in Fig. 7(b), closely match the reference values (RMSE =0.009=0.009).
@@
Figure 7: Experimental results for the response $x^{\prime}=1-T$, where T is the transmission of a $5.5$ cm-long HCN gas cell at $25$ Torr, centered at $1545.23$ nm. (a) Series of the measured powers $b_{i}$ as a function of the modulation index $\beta/\pi$. (b) Reconstructed spectral values of $x^{\prime}$, (blue squares) and the reference curve (red line) obtained from NIST data. The RMSE of the reconstruction is $0.009$.
TABLE I: Summary of parameter values from simulations and experiments